elimination method
4x+5y=2
8x+10y=4

40x+50y=20
40x+-50y=40
80x=40
x=.5
4(.05)+5y=2
2+5y
y=4/5

Not quite sure if this is correct Can you help me please

Answer by user_dude2008(1862)

4x+5y=2
8x+10y=4

-8x-10y=-4
8x+10y=4

0=0

Infinite number of solutions